if a right triangle has a leg measuring 52 m and a hypotenuse measuring 63 m. What is the length of the other leg?

Answer by checkley77(12844)

A^2+B^2=C^2
52^2+B^2=63^2
2,704+B^2=3,969
B^2=3,969-2,704
B^2=1,265
B=SQRT1,265
B=35.567m. ANS. FOR THE SECOND LEG.